# Extreme Jet Beaming Observed in Neutrino-Associated Blazars

### v1

@@ 4.2 Beamed neutrino emission implications
Figure 5: A visual comparison of the typical (median) blazar geometry for neutrino-coincident blazars (Table 1) and for the complete MOJAVE 1.5JyQC sample (Homan et al., 2021). The angles are shown to scale with gray lines indicating 1 and 5 degree steps for guidance. The visual scaling of the velocity and beaming factors is arbitrary.
@@
Neutrino production at $\sim$ parsec distances from the central black hole is consistent with the likely scenario of $p\gamma$ interactions within the jet. There, the required X-ray target photon field can be produced either by the jet itself (self- or external Compton, Plavin et al., 2021; Kalashev et al., 2023) or by other regions of the AGN, such as the corona (e.g., Inoue et al., 2019; Murase et al., 2020; Sotirov, 2024). The exact origin of these photons remains an open question, together with the specific mechanism accelerating protons to PeV energies.
We note that our findings do not preclude the presence of other, non-beamed, sources of neutrinos. The neutrino–VLBI flux density and neutrino–Doppler factor correlations imply that there is a population of blazars that emit neutrinos, and emit them predominantly along the jet axis. Non-beamed and non-jet-related neutrino production, such as possible for Seyfert-type AGNs like NGC1068 (IceCube Collaboration et al., 2022; Abbasi et al., 2024b; Neronov et al., 2024), may form another population of sources leading to a diverse extragalactic neutrino sky.
Heavily flaring blazars producing neutrino during their flares may potentially form another population of neutrino sources that our analysis is not sensitive to. Earlier studies reported observational evidence for neutrinos coming preferentially during electromagnetic flux increases, especially in the radio band: P20; Allakhverdyan et al. (2024b); Hovatta et al. (2021); Albert et al. (2024); Kouch et al. (2024). Those sources may or may not demonstrate extreme jet beaming and persistently bright radio emission. For example, TXS 0506+056 exhibited barely any activity before the neutrino arrival and the following major radio flare in 2018 – 2022. This likely affects its beaming estimates, see Section 4.1. Meanwhile, PKS 1502+106 is another heavily flaring blazar that was originally selected for its highest temporal neutrino-radio correlation in P23 and later studied as a likely neutrino counterpart (e.g. Oikonomou et al., 2021; Britzen et al., 2021). Unlike TXS 0506+056, this blazar demonstrates an extremely strong beaming in our analysis, see Table 1.
5 Conclusions
@@
The observational connection between jet electromagnetic emission beaming and detected neutrinos requires the neutrino emission to be highly beamed as well. The beaming effects appear to be comparable for electromagnetic and neutrino emission, but quantifying them would require both more detailed modeling (to handle observational biases and the intrinsic diversity of blazars) and robust individual neutrino associations. Neutrino beaming implies that the multi-PeV protons producing them exhibit a strong bulk motion along the jet direction. In typical jet acceleration profiles, this motion is only attained at parsec scales, potentially providing a lower limit on the distance from the central black hole to the neutrino production region.
Evaluating direct jet beaming indicators of neutrino-coincident blazars, such as the Doppler factor, already draws a consistent picture of their jets as extremely relativistic and aligned with the line of sight. These indicators are challenging to measure and can only be determined for a limited subset of blazars with great observational coverage. Directly observable properties such as the parsec-scale emission strength measured by VLBI remain well-suited for larger scale searches for neutrino – blazar associations.
@@ Acknowledgements
We thank Eduardo Ros for comments on the manuscript, Markus Böttcher, Anna Franckowiak, and Emma Kun for productive discussions. A.V.P. is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273). Y.Y.K. is supported by the MuSES project which has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No 101142396). The work of S.V.T. is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations.
This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of data from the Radio Fundamental Catalog (Petrov & Kovalev, 2025, doi:10.25966/dhrk-zh08). This research made use of the data from the MOJAVE database maintained by the MOJAVE team (Lister et al., 2018).
@@ Appendix A Beaming of neutrino-coincident blazars in the full MOJAVE sample
In the main analysis of this paper, Section 4.1, we utilize the uniformly-selected subsample of the full set of blazars observed within the MOJAVE program (1.5JyQC). That subset is purely flux-density limited (Section 2.1), leading to the most robust findings. Still, it is potentially instructive to look at the full MOJAVE blazar sample consisting of 447 sources with measured Doppler boosting (Homan et al., 2021). The comparison is shown in Figure A1. The presence of relatively fainter blazars makes the neutrino-coincident ones stand out even more than in the 1.5JyQC analysis (Figure 4). This effect highlights the importance of population studies, large-sample survey and monitoring programs.

### v2

@@ 4.2 Beamed neutrino emission implications
Figure 4: Schematic illustration comparing the typical geometries for blazars in the MOJAVE 1.5JyQC sample (top; D. C. Homan et al. 2021) and neutrino-coincident blazars (bottom; Table 1). Angles are shown to scale, with gray lines indicating $1\degree$ and $5\degree$ increments for guidance. All quantities depicted are medians within their respective samples. The visual scaling of velocity and beaming factors is arbitrary.
@@
Neutrino production at $\sim$ parsec distances from the central black hole is consistent with the likely scenario of $p\gamma$ interactions within the jet. There, the required X-ray target photon field can be produced either by the jet itself (self- or external Compton, A. V. Plavin et al., 2021; O. Kalashev et al., 2023; D. F. G. Fiorillo et al., 2025) or by other regions of the AGN, such as the corona (e.g., Y. Inoue et al., 2019; K. Murase et al., 2020; S. Sotirov, 2024). The exact origin of these photons remains an open question, together with the specific mechanism accelerating protons to PeV energies.
We note that our findings do not preclude the presence of other, non-beamed, sources of neutrinos. The neutrino–VLBI flux density and neutrino–Doppler factor correlations imply that there is a population of blazars that emit neutrinos, and emit them predominantly along the jet axis. Non-beamed and non-jet-related neutrino production, such as possible for Seyfert-type AGNs like NGC1068 (R. Abbasi et al., 2022, 2024b; A. Neronov et al., 2024), may form another population of sources leading to a diverse extragalactic neutrino sky.
Heavily flaring blazars producing neutrino during their flares may potentially form another population of neutrino sources that our analysis is not sensitive to. Earlier studies reported observational evidence for neutrinos coming preferentially during electromagnetic flux increases, especially in the radio band: P20; V. A. Allakhverdyan et al. (2024b); T. Hovatta et al. (2021); A. Albert et al. (2024); P. M. Kouch et al. (2024). Those sources may or may not demonstrate extreme jet beaming and persistently bright radio emission. For example, TXS 0506+056 exhibited barely any activity before the neutrino arrival and the following major radio flare in 2018 – 2022. This likely affects its beaming estimates, see Section 4.1. Meanwhile, PKS 1502+106 is another heavily flaring blazar that was originally selected for its highest temporal neutrino-radio correlation in P23 and later studied as a likely neutrino counterpart (e.g. F. Oikonomou et al., 2021; S. Britzen et al., 2021). Unlike TXS 0506+056, this blazar demonstrates an extremely strong beaming in our analysis, see Table 1.
5 Summary
@@
The observational connection between the jet electromagnetic emission beaming and detected neutrinos requires the neutrino emission to be highly beamed as well. The beaming effects appear to be comparable for electromagnetic and neutrino emission, but quantifying them would require both more detailed modeling (to handle observational biases and the intrinsic diversity of blazars) and robust individual neutrino associations. Neutrino beaming implies that the multi-PeV protons producing them exhibit a strong bulk motion along the jet direction. In typical jet acceleration profiles, this motion is only attained at parsec scales, potentially providing a lower limit on the distance from the central black hole to the neutrino production region.
Evaluating direct jet beaming indicators of neutrino-coincident blazars, such as the Doppler factor, already draws a consistent picture of their jets as extremely relativistic and aligned with the line of sight. This effect is demonstrated here by population analysis, and complemented by dedicated studies of individual sources. In particular, see Y. Y. Kovalev et al. (2025) that finds and explores the extreme jet beaming in the neutrino-associated TeV blazar with low VLBI apparent speed PKS 1424+240 (R. Abbasi et al., 2022). Note that these beaming indicators are challenging to measure and can only be determined for a limited subset of blazars with great observational coverage. Directly observable properties such as the parsec-scale emission strength measured by VLBI remain well-suited for larger scale searches for neutrino – blazar associations.
@@ Acknowledgements
We thank Eduardo Ros and an anonymous referee for comments on the manuscript, Markus Böttcher, Anna Franckowiak, and Emma Kun for productive discussions. A.V.P. is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273). Y.Y.K. was supported by the MuSES project, which has received funding from the European Union (ERC grant agreement No 101142396). The work of S.V.T. is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The views and opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations.
This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of data from the Radio Fundamental Catalog (L. Y. Petrov & Y. Y. Kovalev, 2025, doi:10.25966/dhrk-zh08). This research made use of the data from the MOJAVE database maintained by the MOJAVE team (M. L. Lister et al., 2018).
Code to reproduce most of the figures from this manuscript can be found at https://github.com/aplavin/astroplots.jl/tree/master/nubeam25.
@@ Appendix A Beaming of neutrino-coincident blazars in the full MOJAVE sample
Figure A1: MOJAVE VLBI images for the nine 1.5JyQC blazars coincident with IceCube neutrinos from both samples of Section 2. We show here averages of multi-decade individual VLBI images, which provide a better dynamic range to trace the overall structure of the jet. These stacked images are available from the MOJAVE website (https://www.cv.nrao.edu/MOJAVE/), see details in A. B. Pushkarev et al. (2017). For each blazar, we show the redshift $z$, the Doppler factor $D$ and the viewing angle $\theta$ estimates. Blazars are ordered by their Doppler factor, same as in Table 1.
In the main analysis of this paper, Section 4.1, we utilize the uniformly-selected subsample of the full set of blazars observed within the MOJAVE program (1.5JyQC). That subset is purely flux-density limited (Section 2.1), leading to the most robust findings. The VLBI images of the nine 1.5JyQC blazars coincident with neutrinos are shown in Figure A1. Still, it is potentially instructive to look at the full MOJAVE blazar sample consisting of 447 sources with measured Doppler boosting (D. C. Homan et al., 2021). The comparison is shown in Figure A2. The presence of relatively fainter blazars makes the neutrino-coincident ones stand out even more than in the 1.5JyQC analysis (Figure 3). This effect highlights the importance of population studies, large-sample survey and monitoring programs.